A school administrator claims that less than 50% of the students of the school are dissatisfied



by the community cafeteria service. Test this claim by using sample data obtained from a survey of



500 students of the school where 54% indicated their dissatisfaction of the community cafeteria



service. Use 𝐠= 0.05.




1) Complete the table.



n X p₀ z p-value



a) 120 21 5%



b) 138 32 7%



c) 200 45 10%



d) 392 102 18%



e) 612 236 20%



f) 100 40 8%



g) 248 51 10%



h) 312 100 12%



1
Expert's answer
2022-06-07T15:31:21-0400

When the sample size is large the sample proportion is normally distributed.



"np=500(0.5)=250>30""nq=500(0.5)=250>30"

With "n=500," the Central Limit Theorem applies.

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p\ge0.50"

"H_a:p<0.50"

This corresponds to a left-tailed test, for which a z-test for one population proportion will be used.

Evidence:

Based on the information provided, the significance level is "\alpha = 0.05

," and the critical value for a left-tailed test is "z_c = -1.6449."

The rejection region for this left-tailed test is "R = \{z: z < -1.6449\}."

The z-statistic is computed as follows:



"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{0.54-0.5}{\sqrt{\dfrac{0.5(1-0.5)}{500}}}=3.5777"

Since it is observed that "z =3.5777 \ge-1.6449= z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=P(Z<3.5777)= 0.999827," and since "p=0.999827>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population proportion "p" is less than 0.50, at the "\alpha = 0.05" significance level.




"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}"

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 n & X & p_0 & z & p-value \\ \hline
 120 & 21 & 0.05 & 6.2828 & 0 \\
 \hdashline
 138 & 32 & 0.07 & 7.4534 & 0\\
 \hdashline
 200 & 45 & 0.10 & 5.8926 & 0\\
 \hdashline
 392 & 102 & 0.18 & 4.1333 &0.000036 \\
 \hdashline
 612 & 236 & 0.20 & 11.4800 & 0\\
 \hdashline
 100 & 40 & 0.08 & 11.7954& 0\\
 \hdashline
 248 & 51 & 0.10 & 5.5457&0 \\
 \hdashline
 312 & 100 & 0.12 &10.8990 &0 \\

\end{array}"